OpenStudy (anonymous):

sin(21 degrees)= 11/x

OpenStudy (anonymous):

How would you solve for x here: \(\large A=\frac{11}{x} \) ???

OpenStudy (anonymous):

Multiply the 11 over to the other side? Then do it like 11•sin(21)=x?

OpenStudy (anonymous):

Not quite.... multiply the x over to the other side and divide the A over to the other....

OpenStudy (anonymous):

So you should end up with: \(\large x=\frac{11}{A} \) so you should have: \(\large x=\frac{11}{sin(21^{\circ })} \)

OpenStudy (anonymous):

30.7?! Okay I got it I think! Thank you

OpenStudy (anonymous):

yep... that's right.